OpenStudy (anonymous):

Taylor Polynomials?? if they give you an x, but dont give you an a..is it assumed to be a 0?

OpenStudy (anonymous):

if a=0,then theyr called maclauren's....generally u cannot take a=0.

OpenStudy (anonymous):

Find the Taylor Polynomials of 3rd order generated by f(x) = cos(x) at x= pi/4..so wat do we plug in for a?

OpenStudy (anonymous):

a=pi/4 !!!

OpenStudy (anonymous):

oo because in another problem, he said A = something...this time he said X, so i wasnt sure..

OpenStudy (anonymous):

see i guess,ur confused with the x=pi/4. they've asked u to find taylor POLYNOMIAl and so the result will be a polynomial in x. so x=pi/4 doesn't mean here u'll be substituting that value for x. if u do u won't get a polynomial,rather a single value. get it ?

OpenStudy (anonymous):

makes alot of sense... thanks a ton!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

and if ur satisfied click 'good answer' so dat others won't waste time looking to solve question already clarified :)